USE DE MOIVRE'S THEOREM TO EVALUATE (SQUARE ROOT OF 3 MINUS 1)^5. EXPRESS YOUR ANSWER IN THE STANDARD FORM A+BI

1
Expert's answer
2022-01-18T13:45:04-0500

De Moivre's theorem states


"(\cos \theta +i\sin \theta)^n=\cos n\theta+i\sin n\theta"


Let "z=\sqrt{3}-i"


"|z|=\sqrt{(\sqrt{3})^2+(-1)^2}=2"

"\tan \theta=\dfrac{-1}{\sqrt{3}}, \theta=-\dfrac{\pi}{6}"

"z=\sqrt{3}-i=2(\cos (-\dfrac{\pi}{6}) +i\sin (-\dfrac{\pi}{6}))"

Therefore


"z^5=(\sqrt{3}-i)^5=(2(\cos (-\dfrac{\pi}{6}) +i\sin (-\dfrac{\pi}{6})))^5"

"=2^5(\cos (-\dfrac{\pi}{6}) +i\sin (-\dfrac{\pi}{6}))^5"

"=32(\cos (-\dfrac{5\pi}{6}) +i\sin (-\dfrac{5\pi}{6}))"

"=32(-\dfrac{\sqrt{3}}{2}-i\dfrac{1}{2})"

"=-16\sqrt{3}-16i"